Supplemental Operations Report March 20th

Supplemental Operations Report 20 Mar 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Outpost water supply

Action taken for non-nominal systems: We drained the water tank and refilled it halfway, then shocked it. Water still suspect, so continuing to fill the tank and shocked it with a half-gallon of bleach this time. We will wait three days, then drain and mop out the bottom, then refill the tank. At that time, we will refill and shock the trailer tanks, then empty and refill them. The treatment will be complete and the water should be good again. In the meantime, we will use water pumped directly into the trailers from the HabCar and drink bottled water. An addition action required from now on of the water hoses is that they be cleaned with bleach wipes each time they are used.

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – not noted % (It’s windy and I took a shower so I’m not going out to read them)

Propane Reading, director tank— not noted %

Propane Reading, intern tank— not noted %

Propane Reading, generator— not noted %

Ethanol Free Gasoline – 1 gallon

Water (Outpost tank) – 450 gallons of non-usable water

Sojourner rover used: yes, currently being used by the crew, who promised to treat her nicely.

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: Still using Percy as well as Soji.

ATV’s Used: (Honda, 350.1, 350.2, 300): Honda and the 350 that runs (I always forget to note which one it is)

Reason for use: maintenance and to be used by the crew

Oil Added? Yes to 350, which was basically empty. The Honda levels are fine, but the oil is very dirty.

ATV Fuel Used: 2 gallons

# Hours the ATVs were Used: 1 hour

Notes on ATVs: Turned over today for use by the crew.

HabCar used and why, where? Yes, to town for garbage and water runs

CrewCar used and why, where? No, it is parked in Grand Junction

Luna used and why, where? Yes, to town and for school field trips

Campus wide inspection, if action taken, what and why? The conduit between the Hab and Outpost was smashed right were the opening is at its most narrow. About a 2-foot section has been destroyed. I will work on fixing it this week. The opening will be closed to vehicles.

General notes and comments: Nothing to report

Summary of general operations: Crew arrived on Friday night to set up a 26 foot dome for their class of 28.

Summary of internet: I’m going to switch the Hab internet to Starlink because it works well and has no data limit. The Hab HughesNet has still not been serviced but is still working. The astronomy internet has been restored and is nominal. The Outpost internet is nominal, but I am currently using Starlink.

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Outpost operations: James left today so intern trailer is empty. We moved the double mattress to the bedroom and both bunks currently have the new twin beds. Water issues continue.

Summary of GreenHab operations: Ogden Preparatory Academy, who was here the second half of the week, cleaned the GreenHab as a service project, floors and surfaces were scrubbed and the 60% shade cloth was added to protect things from sun damage over the summer. The heat was accidently turned on during their work and I discovered it at 5 am Saturday morning. The water tank was emptied for the season, and the drain was cleaned with the remaining water from the tank.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: We were able to manually open the robotic observatory dome roof. Once inside we were able to readjust it to open and close correctly, reactivated the internet, installed the new computer mount and get things operational for Peter to test. Testing showed that one cord is wrapped too tightly around the mount and will need to be adjusted. This will be done as soon as the wind lets us.

Summary of health and safety issues: Nothing to report


Copyright © The Mars Society. All rights reserved. | Main Site